OpenStudy (anonymous):

Which of the following numbers can be classified as terminating decimals? Choose all that apply. A. 5.6792 B. 5.5 C. square root of 2 D. 0.67 E. 0.318901 Please help me im very confused!

OpenStudy (anonymous):

what is it that confuses you

OpenStudy (anonymous):

i could just give you the answer but i need to know how are you confused

OpenStudy (anonymous):

I dont know what a terminating decimal is

OpenStudy (anonymous):

it wont open x:

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

yeah:/

OpenStudy (anonymous):

A terminating decimal means a decimal with a definite end. For example 1/2 = .50. 3/4 = .75 3/8 = .375 The decimals that don't end or called non-terminating decimals.

OpenStudy (anonymous):

So would the answer be ABDE?